OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Enter your answer as an ordered pair. y=8x-4 x=4

OpenStudy (anonymous):

y=8x-4 y=8(4)-4 y=32-4 y=28

OpenStudy (anonymous):

huhh??

OpenStudy (anonymous):

(x,y)=(4,28)